### v2

# supplement.pdf

Abstract
We propose a class of metrological resource states whose quantum Fisher information scales optimally in both system size and noise rate. In these states, qubits are partitioned into sensing groups with relatively large correlations within a group but small correlations between groups. The states are obtainable from local Hamiltonian evolution, and we design a metrologically optimal and efficient measurement protocol utilizing time-reversed dynamics and single-qubit on-site measurements. Using quantum domino dynamics, we also present a protocol free of the time-reversal step that has an estimation error roughly twice the best possible value. Finally, we show that spin squeezed states are also optimal for noisy metrology under general conditions.
Introduction.— Quantum metrology Giovannetti et al. (2011); Degen et al. (2017); Pezzè et al. (2018); Pirandola et al. (2018) is the study of utilizing entanglement and other quantum effects to measure unknown data, such as gravitational waves Caves (1981); Yurke et al. (1986); LIGO Collaboration (2011, 2013), images Le Sage et al. (2013); Lemos et al. (2014); Tsang et al. (2016); Abobeih et al. (2019), magnetic fields Wineland et al. (1992); Bollinger et al. (1996); Leibfried et al. (2004); Taylor et al. (2008); Zhou et al. (2020), and time Rosenband et al. (2008); Appel et al. (2009); Ludlow et al. (2015); Kaubruegger et al. (2021); Marciniak et al. (2022). A typical metrological protocol involves initializing $N$ probes in a particular (possibly entangled) state, letting the value of an unknown quantity $\theta$ be imparted on the state, and estimating $\theta$ via a measurement. The Heisenberg limit (HL) Giovannetti et al. (2006) restricts the best-case estimation precision to scale inversely with the number $N$ of entangled probes, $\delta\theta\sim 1/N$.
The prototypical example of a “metrologically optimal” state is the Greenberger-Horne-Zeilinger (GHZ) state Giovannetti et al. (2006), which is used to attain the HL in estimating a $Z$-axis magnetic field in a noiseless $N$-spin system. Unfortunately, the useful but fragile quantum correlations of this state quickly break down in the presence of noise. Instead of attaining the HL, the optimal scaling in a generically noisy setting is the standard quantum limit (SQL), $\delta\theta\sim 1/\sqrt{N}$ Giovannetti et al. (2006); Demkowicz-Dobrzański et al. (2012); Escher et al. (2011); Demkowicz-Dobrzański and Maccone (2014); Sekatski et al. (2017); Demkowicz-Dobrzański et al. (2017); Zhou et al. (2018).
For quantum metrology to continue to bear fruit, it is important to quantify the performance of noise-robust metrological protocols not only in terms of the number of probes $N$ Giovannetti et al. (2006); Demkowicz-Dobrzański et al. (2012); Escher et al. (2011); Demkowicz-Dobrzański and Maccone (2014); Sekatski et al. (2017); Demkowicz-Dobrzański et al. (2017); Zhou et al. (2018), but also in terms of the noise rate $p$ of each probe. A more precise calculation for several models with non-negligible noise yields a precision of $\delta\theta\gtrsim\sqrt{p/N}$  Huelga et al. (1997); Demkowicz-Dobrzański et al. (2012); Escher et al. (2011); Demkowicz-Dobrzański and Maccone (2014); Zhou and Jiang (2021), meaning that reducing the noise rate $p$ can lead to a $\sqrt{p}$ improvement on top of the naive scaling with $1/\sqrt{N}$. While the improvement can be significant for a small $p$, it was rarely explored previously what types of quantum states and measurement protocols can achieve it. Numerical evidence Jarzyna and Demkowicz-Dobrzański (2013); Chabuda et al. (2020) suggests that matrix product states with low bond dimension may be optimal for strong noise. However, these results are based on specific examples, without a general theoretical framework, and the underlying mechanisms behind this behavior remain unclear.
Figure 1: Four families of metrological optimal states and their corresponding preparation and measurement protocols. The first three satisfy the small correlation condition Eq. (6), and each row is a special case of the previous row as indicated by $\subseteq$. Each nested subset admits a more efficient scheme for state preparation and measurement than the previous one. Spin-squeezed states, described in the fourth row, violate Eq. (6) and do not belong to the previous families.
In this Letter, we pioneer the study of quantum estimation limits as functions of both $N$ and $p$. We successfully identify both criteria on and examples of protocols whose variance scales optimally  [^1] We say a family of quantum states $\rho$ as functions of $N$ and $p$ is optimal, if and only if for any $N$ and $p$, there exists a universal constant $c$ such that the QFI ${\mathcal{F}}(\rho_{\theta})\geq c\min\{N/p,N^{2}\}$. with $N$ and $p$. We also construct efficient state preparation and measurement protocols of families of metrologically optimal states. See Fig. 1 for a summary.
We first show that a careful tuning of the quantum correlation length allows for the achievement of optimal scalings. The extensive entanglement of the GHZ state is optimal only in the relatively noiseless setting of $pN\sim$ constant Hayashi et al. (2022), while unentangled states do not yield any $p$-dependent precision improvement. Interpolating between the two edge cases, we show that a “Goldilocks” amount of quantum correlation—quantified by a correlation length $M\sim 1/p$—is optimal. Then we construct states that satisfy our criteria for optimality using local-Hamiltonian evolution and present a provably efficient measurement protocol utilizing time-reversed evolution (similar to the Loschmidt echo Macrì et al. (2016)) and single-qubit measurements. We showcase our results by an example of optimal states generated by analytically solvable quantum domino dynamics Lee and Khitrin (2005); Yoshinaga et al. (2021), where local measurements without time reversal suffice for near-optimal performance. Finally, we establish the optimality of a different class of states that may be long-range correlated: spin squeezed states.
Setting and example.—In this work, we will state our optimality conditions for sensing a $Z$-axis magnetic field under arbitrary noise, with a straightforward extension to any single-qudit Hamiltonian rotations presented in SM .
We use dephasing noise as an illuminating warm-up example. Let $N$ qubits, labeled by $i\in\Lambda$ with Pauli matrices $X_{i},Y_{i},Z_{i}$, be prepared in a state $\rho$, with each qubit $i$ undergoing the channel
${\mathcal{D}}_{\theta,p}(\rho_{i})=(1-p)\mathrm{e}^{-\mathrm{i}\theta Z_{i}}\rho_{i}\mathrm{e}^{\mathrm{i}\theta Z_{i}}+pZ_{i}\mathrm{e}^{-\mathrm{i}\theta Z_{i}}\rho_{i}\mathrm{e}^{\mathrm{i}\theta Z_{i}}Z_{i}$ (1)
for $20<p<1/$ 2. For simplicity, we assume $p>1/N$ (otherwise the noise is asymptotically negligible and the HL is achievable). One then measures the final state $\rho_{\theta}={\mathcal{D}}_{\theta,p}^{\otimes N}(\rho)$ to estimate $\theta$.
The estimation precision can be bounded by the state’s quantum Fisher information (QFI), $\delta\theta\geq 1/\sqrt{N_{\rm e}{\mathcal{F}}(\rho_{\theta})}$, via the quantum Cramér-Rao bound Helstrom (1976); Holevo (2011); Braunstein and Caves (1994); Barndorff-Nielsen and Gill (2000); Gill and Massar (2000); Paris (2009), saturable as the number of repeated experiments $N_{\rm e}\rightarrow\infty$. The QFI for the above example is bounded by Huelga et al. (1997); Escher et al. (2011); Demkowicz-Dobrzański et al. (2012); Zhou and Jiang (2021)
${\mathcal{F}}(\rho_{\theta})\leq 4N/[1-(1-2p)^{2}]\sim N/p,$ (2)
yielding the aforementioned scaling of $\delta\theta\sim\sqrt{p/N}$  [^2] We use $f\sim g,f\gtrsim g$ and $f\lesssim g$ to denote $c_{1}g\leq f\leq c_{2}g,f\geq c_{3}g$ and $f\leq c_{4}g$, respectively, where $c_{1,2,3,4}$ are positive constants., which also applies to general noise channels. Note that Eq. (2) cannot be attained by separable states, for which ${\mathcal{F}}(\rho_{\theta})\leq 4N$, because the QFI of any single-qubit state is upper bounded by $4$. This QFI bound cannot be surpassed with arbitrary state preparation and measurement protocols, including quantum error correction Sekatski et al. (2017); Zhou et al. (2018); Demkowicz-Dobrzański et al. (2017).
[…]
To saturate the QFI upper bound (up to a constant factor), we can partition the $N$-qubit system into $N/M$ groups of $M$ qubits each (with $N/M$ assumed to be an integer for simplicity), tuning the groups’ size inversely with the noise rate, and repeating the GHZ protocol within each group.
Specifically, we consider the quantum parity state, $\rho=\bigotimes_{I}\ket{\mathrm{GHZ}}_{I}\bra{\mathrm{GHZ}}_{I}$ with |$\ket{\mathrm{GHZ}}_{I}=\frac{1}{\sqrt{2}}(\ket{0}^{\otimes M}+\ket{1}^{\otimes M})$ (an instance of the quantum parity code [Uncaptioned image] Shor (1995); Knill et al. (2000); Ralph et al. (2005)). To find a suitable choice of $M$, we first note that in the noiseless limit $p=0$, each group $I$ gains a phase $M\theta$ independently that corresponds to a QFI $\sim M^{2}$, so the QFI of the entire state, equal to the sum of individual QFIs, scales as $(N/M)M^{2}=NM$. In addition, this scaling is robust against noise if $pM$ is upper bounded by a constant, because then dephasing noise will only cause the individual QFI to reduce by a constant factor Huelga et al. (1997). Therefore, Eq. (2) is saturated by tuning $M\sim 1/p$. Replacing the GHZ state in each group by another state that achieves the HL within the group, such as a spin-squeezed state Kitagawa and Ueda (1993); Ma et al. (2011); Ulam-Orgikh and Kitagawa (2001); Brask et al. (2015); Sørensen and Mølmer (2001), does not change that arguments above and yields another optimal state.
The state above has a tensor-product structure w.r.t. the partitioning into groups. This may be too strict of a condition for practical purposes, e.g., when nearest neighbor interactions in the quantum sensor cannot be turned off. Therefore, it is crucial to understand the robustness of this partitioning strategy in the presence of correlations among the groups. This is a nontrivial task since the QFI of noisy mixed states is difficult to analyze and has no obvious relation with usual correlation measures Tóth and Apellaniz (2014). We solve this problem below by showing that any state with sufficiently large intra-group and sufficiently small inter-group correlations achieves the optimal QFI scaling $\sim N/p$. Moreover, our results throughout hold for general noise channels: The final state is $\rho_{\theta}=\mathcal{N}_{\theta,p}^{\otimes N}(\rho)$, where $\mathcal{N}_{\theta,p}$ is a general single-qubit noisy channel depending on $\theta$, and the noise parameter $p$ quantifies the diamond distance between noiseless ($\mathcal{N}_{\theta,0}$) and noisy ($\mathcal{N}_{\theta,p}$) evolutions (see (SM, , Sec. 5.1) for precise definitions).
Sufficient condition for optimality.— Before stating our general sufficient condition (see SM for the proof), we first introduce some notation. For any state $\sigma$ and operators $A,B$, define the expectation $\left\langle A\right\rangle_{\sigma}:=\mathrm{tr}\left(\sigma A\right)$, the correlation function
$C_{\sigma}(A,B):=\frac{1}{2}\left\langle AB+BA\right\rangle_{\sigma}-\left\langle A\right\rangle_{\sigma}\left\langle B\right\rangle_{\sigma},$ (3)
[…]
Theorem 1.
For any given $p$, choose any $M<N$ with
$M\leq c_{\rm no}/p,$ (4)
where $c_{\rm no}$ is a positive constant, and partition the qubits into $N/M$ groups as above. Given any state $\rho$, suppose for any group $I$, its QFI in the noiseless limit is large,
${\mathcal{F}}\left(\rho^{(0)}_{I,\theta}\right)\sim M^{2},$ (5)
and its correlations with other groups are small
$\sum_{J:J\notin\mathcal{S}_{I}}C_{\rho}(I;J)\lesssim M^{-2},$ (6)
where $\mathcal{S}_{I}$ is some set of groups containing a finite (independent of $N,p,M$) number of elements (including $I$) that may have large correlation with $I$. Then there exists an observable ${\mathcal{O}}=\sum_{I}{\mathcal{O}}_{I}$ such that
${\mathcal{F}}(\rho_{\theta})\geq\frac{1}{\left\langle\Delta{\mathcal{O}}^{2}\right\rangle_{\rho_{\theta}}}\left(\partial_{\theta}\left\langle{\mathcal{O}}\right\rangle_{\rho_{\theta}}\right)^{2}\sim NM.$ (7)
Note that $\left\langle\Delta{\mathcal{O}}^{2}\right\rangle_{\rho_{\theta}}$ and $\left\langle{\mathcal{O}}\right\rangle_{\rho_{\theta}}$ represent its variance and expectation value w.r.t. $\rho_{\theta}$, respectively. See, e.g., Ref. Pezzé and Smerzi (2009) for the lower bound on QFI in Eq. (7). ${\mathcal{O}}$ here give rise to an optimal observable to measure, which is a sum of individual observables ${\mathcal{O}}_{I}$ on each group. It implies individual measurements on each group suffice to achieve the optimal QFI, which may still be challenging to implement in cases when the group size $M$ is large.
Theorem 1 implies the optimal QFI scaling (Eq. (2)) is achieved for a family of states when $M\sim 1/p$ and two conditions in Eq. (5) and Eq. (6) are satisfied. They generalize the quantum parity state example:
  1. (i)
    Large intra-group correlation: Eq. (5) requires that each group has the asymptotically optimal QFI $\sim M^{2}$  [^3] In SM , we show that when $\rho$ is pure, the QFI of each RDM can be interpreted as certain correlation functions of $\rho$.. (For the quantum parity state, the RDM of each group is a $M$-qubit GHZ state, satisfying this condition.)
  2. (ii)
    Small inter-group correlation: Eq. (6) requires each group to have a large correlation with at most a finite number of other groups. (For the quantum parity state, each group has zero correlations with others.)
To prove Theorem 1, we first use the large intra-group correlation condition to prove that the QFI is $\sim NM$ when $p=0$, similar to the quantum parity state. Next, we use the small inter-group correlation condition to prove, when $M$ is sufficiently small, i.e., $M\sim 1/p$, that the QFI is only negligibly perturbed by noise SM . We also adopt similar proof ideas for all other theorems below, where we approach the problem by first calculating the noiseless case and then show negligible perturbation under noise. As a result, our theorems apply for general Hamiltonian and noise models. As we will see below, our general criteria for optimal noisy metrology can be satisfied by a large family of many-body states, where efficient preparation and measurement protocols exist.
Locally generated optimal states.— We show Theorem 1 can be naturally satisfied by states generated by local Hamiltonian evolution. Here, local Hamiltonians are sums of geometrically local terms supported on a constant-size sphere on a lattice.
Suppose the qubits lie on vertices of a $d$-dimensional square lattice, start from |$\ket{\bm{0}}:=\ket{0}^{\otimes N}$. We consider states |$\ket{\psi}=U\ket{\bm{0}}$, where unitaries $U$ are generated by a (possibly time-dependent) local Hamiltonian for time $T$. Such evolution is quasi-causal because it satisfies the following Lieb–Robinson bound Lieb and Robinson (1972); Chen et al. (2023): if an operator $A$ acts within set $S\subseteq\Lambda$, then, in the Heisenberg picture,
$A(T):=U^{\dagger}AU\approx\widetilde{A}_{\widetilde{S}},$ (8)
i.e., $A(T)$ can be approximated (up to small error) by an operator $\widetilde{A}_{\widetilde{S}}$. It is strictly supported in the “light cone” region
$\widetilde{S}:=\{i\in\Lambda:\mathsf{d}(i,S)\leq T\},$ (9)
This region includes all vertices within a distance $T$ from $S$. Note that here $T$ is rescaled such that the Lieb–Robinson velocity Chen et al. (2023) is set to $1$, $\mathsf{d}(\cdot,\cdot)$ is the distance function with $\mathsf{d}(i,S)=\min_{j\in S}\mathsf{d}(i,j)$, and we will assume Eq. (8) is exact from now on, with the rigorous treatment given elsewhere SM .
[…]
To satisfy the condition Eq. (5), we first choose $T\sim L$ so that qubits within each group are sufficiently correlated so as to satisfy Eq. (5) Chu et al. (2023). Then Eq. (5) is equivalent to the requirement of a large intra-group $ZZ$ correlation 33footnotemark: 3, i.e.,
$C_{\rho}(Z_{I},Z_{I})\geq c_{Z}L^{d}T^{d},$ (10)
where $Z_{I}=\sum_{i\in I}Z_{i}$ and $c_{Z}\sim 1$.
We will give an example |$\ket{\psi}=U\ket{0}$ that satisfies Eq. (10) later. For now, we consider any optimal state that is efficiently preparable by local Hamiltonian evolution, and for which we propose an efficient measurement protocol below.
On-site measurement scheme with time-reversal.— For locally generated states, we propose an on-site measurement scheme by assuming the ability to time-reverse the evolution $U$ (e.g., by changing the sign of the generating Hamiltonian).
After sensing, one applies a unitary $U_{\rm rev}:=U^{\dagger}\mathrm{e}^{\mathrm{i}\theta_{\rm pr}Z}$ ($Z:=\sum_{i}Z_{i}$) to the state $\rho_{\theta}$, and then measures in the computational basis before post-processing the data. Here, $\theta_{\rm pr}$ represents prior knowledge of $\theta$, such that $\tilde{\theta}:=\theta-\theta_{\rm pr}$ satisfies $|\tilde{\theta}|\sim L^{-d}$  [^4] The requirement $|\tilde{\theta}|\lesssim L^{-d}$ is a common assumption in metrology because one can always gain the prior knowledge on the “digits” of $\theta$ one by one with negligible overhead until the requirement is satisfied Higgins et al. (2009); Kimmel et al. (2015); Belliardo and Giovannetti (2020). We also assume $|\tilde{\theta}|\gtrsim L^{-d}$, achievable by shifting $\theta_{\rm pr}$ if necessary, so that the slope Eq. (27) is large.. The idea comes from the Loschmidt echo protocol working in the noiseless case Macrì et al. (2016), where one processes the measurement data by simply projecting the state to |$\ket{\bm{0}}$. The probability is
$\mathsf{P}_{\rm echo}:=\left\lvert\bra{\bm{0}}U^{\dagger}\mathrm{e}^{-\mathrm{i}\tilde{\theta}Z}U\ket{\bm{0}}\right\rvert^{2}\!=1-\frac{\tilde{\theta}^{2}}{4}\!\mathcal{F}(\rho_{\theta})+\mathrm{O}(\tilde{\theta}^{4}).$ (11)
The projection is optimal because the corresponding classical Fisher information $\frac{1}{\Delta\mathsf{P}_{\rm echo}^{2}}\left(\partial_{\theta}\mathsf{P}_{\rm echo}\right)^{2}\approx\mathcal{F}(\rho_{\theta})$ at small $\tilde{\theta}$, where $\Delta\mathsf{P}_{\rm echo}^{2}=\mathsf{P}_{\rm echo}(1-\mathsf{P}_{\rm echo})$ is the variance of a Bernoulli random variable with probability $\mathsf{P}_{\rm echo}$.
However, the Loschmidt echo protocol above through measuring the observable |$\ket{\bm{0}}\bra{\bm{0}}$ | no longer works under noise. To guarantee robustness against noise, we need to post-process the data from computational-basis measurement in a different way. Let $P_{\widetilde{I}}:=\bigotimes_{i\in\widetilde{I}}\ket{0}_{i}\bra{0}$ | be projections onto small regions $\widetilde{I}$ (defined via Eq. (9)), we measure instead the overall observable $\left\langle{\mathcal{O}}_{\rm rev}\right\rangle_{\rho_{\theta}}$ with
${\mathcal{O}}_{\rm rev}:=\sum_{I}\widetilde{P}_{I}:=\sum_{I}U_{\rm rev}^{\dagger}P_{\widetilde{I}}U_{\rm rev}.$ (12)
Note that we often omit the identity operator on the rest of the system for simplicity, e.g., here $P_{\widetilde{I}}$ means $P_{\widetilde{I}}\otimes\mathrm{id}$. This time-reversal on-site measurement scheme is asymptotically optimal.
Theorem 2.
For any $p$, choose $L=M^{1/d}$ such that Eq. (4) holds for some constant $c_{\rm no}$ (which is not necessarily the same constant in Theorem 1). For any locally generated state $\rho$ satisfying Eq. (10), there exist constants $C>c>1$ determined by $c_{Z}$ and $d$, such that for any $C\geq L/T\geq c$,
$\frac{1}{\left\langle\Delta{\mathcal{O}}^{2}_{\rm rev}\right\rangle_{\rho_{\theta}}}\left(\partial_{\theta}\left\langle{\mathcal{O}}_{\rm rev}\right\rangle_{\rho_{\theta}}\right)^{2}\sim NL^{d}.$ (13)
Theorem 2 implies locally generated optimal states, combined with the time-reversal measurement protocol, achieve the optimal QFI when choosing $M\sim 1/p$. An additional advantage of the time-reversal measurement protocol is the observable ${\mathcal{O}}_{\rm rev}$ is independent of $p$, in contrast to ${\mathcal{O}}$ in Theorem 1 that may be a function of $p$. A proof sketch can be found in End Matter, with a full proof in SM .
Quantum domino dynamics with on-site measurement.— We have shown that any state generated by local dynamics with correlations of the form Eq. (10) is metrologically optimal when $L^{d}\sim 1/p$. Here we present an explicit example of this framework using quantum domino dynamics Lee and Khitrin (2005); Yoshinaga et al. (2021), where an on-site measurement protocol without time-reversal is sufficient to achieve the optimal QFI.
Consider the $3$-local spin-chain Hamiltonian that flips a spin only if its two neighbors are anti-aligned,
$H_{\rm do}=\frac{1}{2}\sum_{i=2}^{N-1}X_{i}(1-Z_{i-1}Z_{i+1}),$ (14)
which can be engineered as the effective Hamiltonian for the transerve-field Ising model at weak $X$ field Lee and Khitrin (2005). The vacuum |$\ket{\bm{0}}$ is an eigenstate of $H_{\rm do}$, and, more generally, the number of domain walls (DWs)—transitions from |$\ket{0}$ to |$\ket{1}$ or vice versa—is conserved. If the system starts from the initial state |$\ket{1_{1}}:=\ket{10\cdots 0}$, it only traverses the state space with one DW $\mathrm{e}^{-\mathrm{i}TH_{\rm do}}\ket{1_{1}}=\sum_{i=1}^{N}a_{i}\ket{1_{i}}$, where |$\ket{1_{i}}$ is the state with $1$ on the first $i$ qubits and 00 on the rest and the coefficients $a_{i}$ are determined by $T$. This single-particle hopping process of the DW justifies the name “domino”. Before $vT\gtrsim N$ (i.e., when the DW propagates to the right), $a_{i}$ is centered around $i=vT$ with a velocity $v$. Then, an initial state |$\ket{+}_{1}\otimes|\bm{0}\rangle$ would evolve to
|$\ket{\psi_{\rm do}}:=\mathrm{e}^{-\mathrm{i}TH_{\rm do}}\ket{+}_{1}\otimes|\bm{0}\rangle=\frac{1}{\sqrt{2}}\left(\ket{\bm{0}}+\sum_{i=1}^{N}a_{i}\ket{1_{i}}\right),$ (15)
which is roughly a GHZ-like state on the left $\sim vT$ qubits (with other qubits in state |$\ket{0}$), useful for metrology Yoshinaga et al. (2021).
Figure 2: Optimal noisy metrology for $N=200$ qubits under dephasing, using partitioned domino state Eq. (16) generated from quantum domino dynamics, and on-site measurements ${\mathcal{O}}_{\rm do}$ of the form Eq. (20). The $y$-axis quantifies $(\delta\theta)^{-2}/N$ for local estimation at $\theta=0$. For each solid line, the spin chain is divided to $N/L$ groups each of size $L$, and the generating time $T\sim L$ is suitably chosen (see SM for details). For any noise probability $p\ll 1$ in Eq. (1), by choosing suitable $T$ and $L$, the performance approaches the bound in Eq. (2) (pink dashed line) up to a factor $\approx 2$.
Adapting the domino dynamics to our partition framework, we propose to evolve the initial state containing one $\ket{+}$ in every $L$ qubits:
|$\ket{\Psi_{\rm do}}:=\mathrm{e}^{-\mathrm{i}TH_{\rm do}}\left(\ket{+}_{1}\otimes\ket{+}_{L+1}\otimes\cdots\right)\otimes|\bm{0}\rangle_{\rm rest}.$ (16)
This final state, which we call the partitioned domino state, satisfies Eq. (10) for $T\lesssim L$, because the two DWs of each $\ket{+}$ roughly propagate in opposite directions with constant velocity. We numerically study noisy metrology using |$\ket{\Psi_{\rm do}}$ on a chain of $200$ qubits, and Fig. 2 shows the results. Remarkably, this protocol is not only optimal asymptotically, but also quantitatively. By adjusting $L$ for different $p$, the precision $(\delta\theta)^{2}$ is only roughly twice larger than the known lower bound Eq. (2).
Besides having the optimal metrological power, the domino-state protocol requires only on-site measurements without any time-reversed evolution. To show this, we focus first on the simplest one-DW case with state Eq. (15) and prove the following in SM :
Theorem 3.
In a rotated on-site basis (i.e., redefining states |$\ket{1}_{i}$ with extra phases) such that $\mathrm{Arg}(a_{i})=-\frac{\pi}{2}(1+i)-2\theta_{\rm pr}i$, define the measurement observable
${\mathcal{O}}_{\rm do}:=\sum_{vT/2\leq i\leq 2vT}|a_{i}|\mathbf{Y}_{1,i},$ (17)
where $\mathbf{Y}_{1,i}:=\bigotimes_{j=1}^{i}Y_{j}$. If
$\sum_{vT/2\leq i\leq 2vT}|a_{i}|^{2}\geq c_{\rm do}\sim 1,$ (18)
then
$\frac{1}{\left\langle\Delta{\mathcal{O}}^{2}_{\rm do}\right\rangle_{\rho_{\theta}}}\left(\partial_{\theta}\left\langle{\mathcal{O}}_{\rm do}\right\rangle_{\rho_{\theta}}\right)^{2}\sim T^{2},$ (19)
if noise $p\leq c_{\rm do}/(vT)$ for a constant $c_{\rm do}$.
Eq. (17) can be realized by $Y$-basis on-site measurements with efficient classical post-processing. The condition Eq. (18) means that $|a_{i}|^{2}$, as a probability distribution, is concentrated around $vT$, which holds before $vT\gtrsim N$.
The idea to choose Eq. (17) as the observable is as follows. |$\ket{\psi_{\rm do}}$ is roughly a superposition of GHZ states |$\ket{\mathrm{GHZ}}_{1\cdots i}\otimes\ket{\bm{0}}$ of different size $i$. For each GHZ state, $\mathbf{Y}_{1,i}$ is an optimal observable to sense the $Z$-axis magnetic field. However, each such observable does not have large sensitivity, i.e., $\partial_{\theta}\left\langle\mathbf{Y}_{1,i}\right\rangle\sim|a_{i}|i\ll i$, because the DW wave packet spreads. To overcome this, we use a suitable linear combination of them, namely, ${\mathcal{O}}_{\rm do}$, to accumulate the small derivatives into $\partial_{\theta}{\mathcal{O}}_{\rm do}\sim\sum_{i}|a_{i}|^{2}\sim 1$. It turns out this linear combination does not lead to large variance of ${\mathcal{O}}_{\rm do}$ due to the lack of interference among the GHZ states in |$\ket{\psi_{\rm do}}$, making ${\mathcal{O}}_{\rm do}$ a suitable observable.
Generalizing the idea, the observable we use in Fig. 2 for state Eq. (16) is of the form
$\sum_{\text{integer }I}\,\sum_{1-L/2\leq i\leq j\leq L/2}b_{i,j}\mathbf{Y}_{IL+i,IL+j}~{},$ (20)
where the real coefficients $b_{i,j}$ come from simulating the two-DW subspace for the group $I=0$, which only requires polynomial (in $L$) resource of classical computation. See SM for details.
Spin-squeezing.— We have established a sufficient condition for metrological optimality. Here we show that this small-correlation condition (Eq. (6)) is not necessary: spin-squeezed states (SSS) Kitagawa and Ueda (1993); Ma et al. (2011) do not satisfy it in general, but are optimal by the following theorem proven in SM , which generalizes special SSSs and noise analyzed in Ulam-Orgikh and Kitagawa (2001); Zhou and Jiang (2021); Brask et al. (2015).
Theorem 4.
Suppose state $\rho$ satisfies
$\left\langle X\right\rangle_{\rho}$ $=\Theta(N),\quad\left\langle\Delta Y^{2}\right\rangle_{\rho}=N\xi^{2},$ (21)
$\left\langle\Delta S^{2}\right\rangle_{\rho}$ $=\mathrm{O}(N/\xi^{2}),\quad\forall S:=\textstyle\sum_{i}S_{i}\text{ with }\left\lVert S_{i}\right\rVert=1$ (22)
where $X=\sum_{i}X_{i}$ etc and $S_{i}$ denotes an operator on $i$. Its metrological power is robust in noisy metrology with $p=\mathrm{O}(\xi^{2})$:
$\mathcal{F}(\rho_{\theta})\geq\frac{1}{\left\langle\Delta Y^{2}\right\rangle_{\rho_{\theta}}}\left(\partial_{\theta}\left\langle Y\right\rangle_{\rho_{\theta}}\right)^{2}\Big{|}_{\theta=0}=\Theta(N/\xi^{2}).$ (23)
Eq. (21) is a standard definition of SSS: $\rho$ is polarized in the $X$ direction and squeezed by a factor $\xi\ll 1$ in the $Y$ direction. For example, extremely squeezed states with $\xi=\Theta(1/\sqrt{N})$ exist in the permutation-symmetric subspace of the $N$ qubits, for which any two groups of qubits have a constant correlation $C_{\rho}(I;J)=\Theta(1)$, violating Eq. (6). Explanations on the violation of Eq. (6) for any sufficiently small $\xi$ can be found in SM . Although beyond the scope of Theorem 1, such SSSs are optimal by Theorem 4, as long as its “polarization” $S$ along any direction in Eq. (22) is at most stretched by a factor $\mathrm{O}(1/\xi)$, the minimal amount given by the uncertainty principle (as one can see by choosing $S=Z$). The intuition here is that under general noise, the metrological observable $Y$ may mix with “orthogonal” directions $S$ and thus increase its variance; Eq. (22) ensures this mixing is weak enough. The condition Eq. (22) suggests the general usefulness of correlation bounds for noisy metrology.
Eq. (23) shows that on-site measurements suffice for noisy metrology with SSS. However, it is not straightforward to prepare SSSs from natural (i.e., time-independent) local dynamics: The squeezing parameter $\xi$ has a limited scaling with $N$ in recent proposals Block et al. (2024); Comparin et al. (2022); Roscilde et al. (2024), and it is argued that the 1d case can only generate $\xi\sim 1$ independent of $N$. This is in sharp contrast with the preparable domino state above, which has optimal robustness against noise and on-site measurement schemes. Nevertheless, it is unclear whether the domino dynamics can be generalized to higher dimensions. The small-correlation states and SSSs are thus complementary classes of metrological resources, with each offering advantages depending on the context.
Discussion.— In this work, we prove sufficient conditions for optimal noisy metrology and illustrate them with several examples, from simple quantum parity states to quantum domino dynamics and spin-squeezed states.
Beyond state optimality, we also address the problem of efficient measurement schemes for metrology. In particular, the domino example suggests that on-site measurements apply to wider sensing scenarios beyond GHZ: Although an on-site measurement scheme that exactly saturates the QFI Liu et al. (2023) may not exist, the QFI can still be saturated up to a moderate prefactor.
Acknowledgements.— C.Y. was supported by the Department of Energy under Quantum Pathfinder Grant DE-SC0024324. S.Z. acknowledges funding provided by Perimeter Institute for Theoretical Physics, a research institute supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities. VVA thanks Ryhor Kandratsenia and Olga Albert for providing daycare support throughout this work.